OpenStudy (anonymous):

Find the equation of a line which passes through point (1, 7) and is perpendicular to the line formed by the equation -2x + -10y = -9. Express it in slope-intercept form.

OpenStudy (curry):

i already told you y=-5x+12

OpenStudy (curry):

dont forgot medal

OpenStudy (anonymous):

its not 12 its either 2 or -2

OpenStudy (radar):

y=5x+2

OpenStudy (radar):

The line that is given has a slope of -2/10 -10y=2x-9 y=(2/-10)x+9/10 You are interested only in the slope 2/-10 because a line that is perpendicular will have a slope that is a negative reciprocal of that slope or the slope of the perpendicular line will be 5 m=5. since you have a point (given) and you have calculated the slope, you can now write the equation of the line that is perpendicular. y-7 5= ------ cross multiply x-1 5x-5 = y-7 y=5x -5-(-7) y=5x + 2 (answer)